OpenStudy (anonymous):

the equation ans the solution of In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive

OpenStudy (anonymous):

x + y = 221 x - y = 53 solve for x and y

OpenStudy (anonymous):

x+(x+53)=221 2x+53=221 2x=168 x=84. So 84 and 137